OpenStudy (anonymous):

the area of a square boxing ring is 436 square feet.what is the approximate length, in feet, of one side of the boxing ring

OpenStudy (unklerhaukus):

do you know a formula for the area of a square in terms of side lengths?

OpenStudy (anonymous):

Yes

OpenStudy (unklerhaukus):

A=

OpenStudy (unklerhaukus):

...

OpenStudy (anonymous):

L x L = x ^ 2

OpenStudy (unklerhaukus):

yeah, so the question says that, that area equals 436 square feet, so what does that make each side equal?

OpenStudy (anonymous):

I'm not sure it can either be 4,21,54,109

OpenStudy (unklerhaukus):

\[A=L\times L=L^2=436\] \[L=\sqrt{436}\]

OpenStudy (anonymous):

Soo is it 4 or 21

OpenStudy (anonymous):

21

OpenStudy (anonymous):

\[\sqrt{438} = 20.880613017821100359515508045096... \] \[\sqrt{436} \approx 21\] Because 20.8 rounded is 21